OpenStudy (anonymous):

Solve: x^2+8x-8=0 A. {-4, pos./neg. sq. root of 6} B. {4, pos./neg. sq. root of 12} C. {-4, 2 times pos./neg. sq. root of 12} D. {-4, 2 times pos./neg. sq. root of 6}

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@mathmate

OpenStudy (irishboy123):

are you learning how to complete a square? \(x^2+8x-8 = 0\) \( (x+4)^2-16-8 = 0\) ???? or to plug numbers into this (same thing) \( \frac{- b \pm \sqrt{b^2 - 4ac}}{2a}\) ?? or to draw curves?

OpenStudy (anonymous):

it just says to solve it and I don't know how. it's probably completing the square.

OpenStudy (irishboy123):

ok!, well here you go: https://www.mathsisfun.com/algebra/completing-square.html suggest you have a crack at this and report back good luck :p

OpenStudy (mathmate):

@yayah2001 The answer choices are all wrong. The way they are written, there are 3 roots, which is impossible for a quadratic equation. You may want to check if there is a typo, perhaps the comma (,) should be a plus sign or something else.

OpenStudy (anonymous):

I got the answer. thanks for the help!

OpenStudy (irishboy123):

the answer *is* in there :p

OpenStudy (mathmate):

@IrishBoy123 I am not worried about the answer. The OP should understand that there is a difference between {-4, 2 times \(\pm\sqrt{6}\)} which implies 3 roots, and {-4+ 2 times \(\pm\sqrt{6}\)} which means there are two roots. He may have chosen the right letter choice, but real life does not exist in multiple choices, sigh!

OpenStudy (anonymous):

by 2 times I meant multiplication

OpenStudy (triciaal):